OpenStudy (anonymous):

40x-900=-25x+2350 what is the solution set

OpenStudy (mayankdevnani):

40x+25x=2350+900 65x=3250

OpenStudy (mayankdevnani):

can you solve it

OpenStudy (anonymous):

x=50

OpenStudy (anonymous):

is that the solution set? how do i write it?

OpenStudy (mayankdevnani):

yaa

OpenStudy (mayankdevnani):

x=50

OpenStudy (anonymous):

it asks me in this form { } so {50,0}?

OpenStudy (mayankdevnani):

yaa

OpenStudy (anonymous):

the average rate of change of f(x)=5x^2-2, from 2 to 4. can you solve that

OpenStudy (mayankdevnani):

post it

OpenStudy (anonymous):

what do yu mean post it

OpenStudy (unklerhaukus):

zero is not a solution

OpenStudy (unklerhaukus):

\[x=\{50\}\]